Prove and explain how to represent 2n - 1 in binary notation
Answer #1
2^n - 1

if n is 1, then 2^1 - 1 = 2 - 1 = 1. it's binary form is 1
if n is 2, then 2^2 - 1 = 4 - 1 = 3. it's binary form is 11
if n is 3, then 2^3 - 1 = 8 - 1 = 7. it's binary form is 111
so, 2^n - 1 in binary form is 111...1 => (n 1's)